OpenStudy (anonymous):

don and ron together weighed 48 pounds. don weighed two pounds more than ron. how much did ron weighed

OpenStudy (anonymous):

i just need to figure how to write the eauations then i could do it myself

OpenStudy (anonymous):

let rons weight be x hence dons weight be x +2 together weight is 48 hence x + x + 2 = 48 2x + 2 - 2 = 48 - 2 2x = 46 x = 23 hence ron = 23 don = 25

OpenStudy (anonymous):

ok thanks